theyadoreshayyy:

If the cos 30° = square root 3 over 2, then the sin 60° = ________. 0, because the angles are complementary one half, because the angles are complementary square root 3 over 2, because the angles are complementary 1, because the angles are complementary

AZ:

Here's a hint \(\cos x = \sin(90^o - x)\)

AZ:

and similarly \(\sin x = \cos(90^o -x)\)

AZ:

But look at that first identity I gave you. x is 30 They told you \(\cos 30 = \dfrac{\sqrt{3}}{2}\) So what is \(\sin 60 = ?\)